OpenStudy (nightmarenight):

The number of calories burned, C, varies directly with the time spent exercising, t. When Dennis walks for 4 hours, he burns 800 calories. Show the direct linear variation in an equation

OpenStudy (cggurumanjunath):

C = K * t 800 = K * 4

OpenStudy (cggurumanjunath):

K IS CONSTANT OF PROPORTIONALITY.

OpenStudy (cggurumanjunath):

800=4*K 800/4=200=K

OpenStudy (cggurumanjunath):

C=200t

OpenStudy (cggurumanjunath):

C=200t is your required equation .